OpenStudy (anonymous):

Subtract the polynomials: (3x3 - 2x + 3) - (2x3 - 2). Enter your answer as a polynomial in descending order below, and use the caret (^) for exponents. For example, you would write as 4x^2.

OpenStudy (anonymous):

..........

OpenStudy (anonymous):

(3x^3 - 2x + 3) - (2x^3 - 2) \[=(3x^3 - 2x + 3-2x^3+2)\] \[=(3-2)x^3-2x+5\] \[=x^3-2x+5\]